OpenStudy (anonymous):

Write the equation of the line in slope-intercept form that passes through the point (3, -2) and has a slope of 2

OpenStudy (haseeb96):

(y-y1)=m(x-x1) m=2 (x1,y1) =(3,-2) just put it in the formula and get the answer

OpenStudy (haseeb96):

show me yur work move fast

OpenStudy (haseeb96):

y-(-2)=2(x-3) y+2 =2x -6 y=2x-6-2 y=2x-8 this is the equation of a line in slope intercept form